OpenStudy (anonymous):

help!!! 5x+7y=-1 -2y+3z=9 7x-z=27

OpenStudy (technodynamic):

x=4 y=-3 z=1

OpenStudy (anonymous):

you used rref ?

OpenStudy (anonymous):

okay..i need to know how to do them...

OpenStudy (anonymous):

wats rref..this is scaring me...

OpenStudy (anonymous):

techno how did you solve it

OpenStudy (technodynamic):

rref? i don't know what he means by that. But the answer will be...x=4 y=-3 and z=1

OpenStudy (anonymous):

how did you solve it

OpenStudy (anonymous):

okay...but how?

OpenStudy (anonymous):

plz dunt say "i dunt noe? i used mental math"

OpenStudy (technodynamic):

First work on an equation to solve, and then you plug in a variable and you plug it back for the other equation.

OpenStudy (anonymous):

but whadda bout in this case..wad do i do first? and hw did u do it so fast?

OpenStudy (technodynamic):

I got an A in it...it's simple.

OpenStudy (anonymous):

ughh..dunt say its simple...i suck at it and i got a C- in alg 2 rite now...

OpenStudy (anonymous):

ok lets use strategy

OpenStudy (technodynamic):

lol ok

OpenStudy (anonymous):

yes plz..i really need a stragety...to do dis problem and the other zillion

OpenStudy (anonymous):

the two main strategies are elimination and substitution

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

substitution might be faster here, dunno. we can do both ways. ok for elimination we want it to look like this first

OpenStudy (anonymous):

go on...

OpenStudy (anonymous):

5x+7y + 0z = -1 0x -2y+3z = 9 7x+0y -z = 27

OpenStudy (anonymous):

okay....

OpenStudy (anonymous):

wat next?

OpenStudy (anonymous):

lets call them equation 1, 2, 3, . the general strategy is , eliminate x from equation 1 and 2 , (using their lcm) then eliminate x from equations 1 and 3 (using lcm and subtracting each other )

OpenStudy (anonymous):

that leaves you with two equations in two unknowns ( y and z )

OpenStudy (anonymous):

ohhh...i was so confused about the missing x in one of the equations...LOL

OpenStudy (anonymous):

so equation 1 and 2 are 5x+7y + 0z = -1 0x -2y+3z = 9 unfortunately we cant eliminate x here though because second equation has a 0 in front of x

OpenStudy (anonymous):

so lets eliminate x from equations 1 and 3 , you follow so far>

OpenStudy (anonymous):

so we can skip the first step

OpenStudy (anonymous):

kk.im following...

OpenStudy (anonymous):

we will eliminate x from equation 1 and 3 : 5x+7y + 0z = -1 7x+0y -z = 27

OpenStudy (anonymous):

the lcm is 35 , so multiply each other by the missing factor

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

(7)(5x+7y + 0z) = (-1)(7) -5(7x+0y -z) = (27)

OpenStudy (anonymous):

leaving with 35x+49y+5z=-7 -35-5y-5z=-135

OpenStudy (anonymous):

(7)(5x+7y + 0z) = (-1)(7) -5(7x+0y -z) = -5 (27)

OpenStudy (anonymous):

i used -5 because then we can just add

OpenStudy (anonymous):

35 x + 49 y + 0z = -7 -35 x + 0y +5xz = -27*5

OpenStudy (anonymous):

me too...

OpenStudy (anonymous):

how is it 5xz thou?

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

isnt it also 5y thou?

OpenStudy (anonymous):

yes typo

OpenStudy (anonymous):

lol..i thought i was wrong...

OpenStudy (anonymous):

35 x + 49 y + 0z = -7 -35 x + 0y +5z = -27*5

OpenStudy (anonymous):

but wait..isnt it also 5y?

OpenStudy (anonymous):

oh wait...nevermind...

OpenStudy (anonymous):

ok i added them and got 49y + 5z = (-7 - 135)

OpenStudy (anonymous):

now wad do i do?

OpenStudy (anonymous):

now you have two equations in y and z

OpenStudy (anonymous):

49 y + 5z = -142 -2y + 3z = 9

OpenStudy (anonymous):

so u eliminate one of the variables wit tat equation and the second one?

OpenStudy (anonymous):

yes, you can now eliminate z or y

OpenStudy (anonymous):

im thinking the z's rite..cuz tay are smaller numbers

OpenStudy (anonymous):

it looks easier to eliminate z because lcm of 5 and 3 is 15

OpenStudy (anonymous):

yea...

OpenStudy (anonymous):

so once you find what y is, you back substitute

OpenStudy (anonymous):

but wait tat cant work....

OpenStudy (phi):

Here's how to do it using elimination. Ugly numbers...but you get the right answer: This is how to do elimination. It's very methodical, so you have less of a chance getting confused. 5x+7y=-1 -2y+3z=9 7x-z=27 rewrite, using only the coefficients (always in the order x y z) 5 7 0 -1 0 -2 3 9 7 0 -1 27 Now make all the numbers below 5 zero, by multiplying the top row by a number and adding to the row. Here row 2 already has a zero. For row 3, multiply the top row by -7/5 and add to row3. (We pick -7/5 because -7/5 times the 5 in the first row= -7 and -7 +7 in the 3rd row is zero) Here -7/5*(5 7 0 -1 )= (-7 -9.8 0 1.4) 5 7 0 -1 <-- leave alone 0 -2 3 9 <-- leave alone 7 0 -1 27 <-- add (-7 -9.8 0 1.4) (comes from -7/5 times 1st row) 5 7 0 -1 0 -2 3 9 0 -9.8 -1 28.4 now use the 2nd row, 2nd number, and make everything below it zero so multiply the 2nd row by -9.8/2 = -4.9. We get -4.9*(0 -2 3 9)= (0 9.8 -14.7 -44.1) 5 7 0 -1 <-- leave alone 0 -2 3 9 <-- leave alone 0 -9.8 -1 28.4 <-- add (0 9.8 -14.7 -44.1) 5 7 0 -1 <-- leave alone 0 -2 3 9 <-- leave alone 0 0 -15.7 -15.7 Now back substitute: the last line means -15.7z= -15.7, and z= 1 the 2nd line means -2y + 3z= 9 with z=1, we find y= -3 Finally, 5x+7y= -1, with y= -3, we have x= 4

OpenStudy (anonymous):

wow...

OpenStudy (anonymous):

wat a long reply...its been an hour...